Roll a dice n times and let X be the number of times you roll 6. Assume that n is large.

(a) Compute the expectation E[X].

(b) Write down an approximation, in terms on n and , of the probability that X di ers from its

expectation by less than 10%.

(c) How large should n be so that the probability in (b) is larger than 0.99?


1
Expert's answer
2021-05-07T14:16:40-0400

(a) Let X"_i" be the number on the face of die


Let X be the number you roll 6-


Therefore X have the random value as-

       "X=\sum_{i=1}^{n} X_i."


By Linearity of Expectation-


  "E[X]=\sum_{i=1}^{n} E[X_i]."


Also "E[X_i]=\sum_{j=1}^6jP[X_i=j]=\sum_{j=1}^6j(\dfrac{1}{6})=\dfrac{1}{6}\times \dfrac{6(7)}{2}=\dfrac{7}{2}"


Also, "\sum_{j=1}^{n}j=\dfrac{n(n+1)}{2}"


Therefore, "E[X]=\dfrac{7}{2}\times\dfrac{n(n+1)}{2}= \dfrac{7n(n+1)}{4}"


(b) The probability that X differs from its expectation by less than 10% is given by-


"=\sum_{i=1}^{n} X.E(X)=\sum_{i=1}^{n}X.\dfrac{7n(n+1)}{4}"


(c) Probability in b is larger than 0.99-


"\dfrac{7n(n+1)}{4}>0.99\\[9pt]



\Rightarrow 7n^2+7n-3.76>0"


on solving we get, "n=-0.38,3.76"


So The value of n must be 4.